How do you do this problem
2m=n-3
3m=2n-9

Answer by ankor@dixie-net.com(22740)

2m = n - 3
3m = 2n - 9
:
In the standard form:
2m - n = -3
3m -2n = -9
:
Multiply the 1st equation by -2 and add the equations
-4m + 2n = 6
3m - 2n = -9
-----------------adding eliminates n
-m = -3
m = 3
:
Find n using the 1st equation
2(3) - n = -3
6 - n = -3
-n = -3 - 6
n = 9
:
Check solutions in the 2nd equation
3(3) - 2(9) =
9 - 18 = -9